OpenStudy (anonymous):

Where are the asymptotes of f(x) = tan(4x − π) from x = 3 pi over 4 to x = 5 pi over 4? is it x = 0, x = pi / 4??

OpenStudy (anonymous):

is that right?

OpenStudy (xapproachesinfinity):

the tanx is undefined when x=pi/2 and -pi/2 those are the asymptotes. can you do that with tan(4x-pi)

OpenStudy (xapproachesinfinity):

when x=0 tan(-pi)=-tan(pi) which is well defined no asymptotes here

OpenStudy (anonymous):

I have no idea. im in a huge hurry can you please just c if the answer above is right?

OpenStudy (xapproachesinfinity):

none of them is an asymptote since tan is defined what are your choices

OpenStudy (anonymous):

x = pi / 4, x = 3pi / 4. x = 0, x = pi / 4 x = pi / 2, x = 3pi / 2 x = 3pi / 8, x = 5pi /8

OpenStudy (xapproachesinfinity):

it is the last answer do you know why? that is what i was trying to explain to you but you said you are in rush lol

OpenStudy (anonymous):

im sorry but thank you so much!

OpenStudy (xapproachesinfinity):

No problem, I wanted to slowly get you to this. you're welcome

OpenStudy (anonymous):

What are the sine, cosine, and tangent of 2 pi over 3 radians? do u by chance know this one

OpenStudy (xapproachesinfinity):